The two sides of a right triangle are 12 m and 13 m. What is the length of the base of the triangle?

Answer by cleomenius(959)

I will assume the 13 is the hypotenuse.
a^2 + 12*2 = 13^2.
a^2 + 144 = 169.
a = 5 m.
Cleomenius